# ExLM: Rethinking the Impact of [MASK] Tokens in Masked Language Models

### v2

@@ I.4 Space Complexity Analysis
However, as shown in Table 14, ExLM with the Sparse DAG experiences some performance degradation. While it still outperforms the version of ExLM without a transition matrix (ExLM w/o Transitions), it falls behind the original ExLM. The reason for this decline is that, in the Sparse DAG, only one expanded state can be selected for each [MASK] token, which prevents the model from capturing longer but possible outputs. For example, in the sentence ”This is [MASK], and I’m very [MASK] to see this,” both ”so good” and ”happy” are reasonable interpretations for the first and second [MASK] tokens, respectively. However, the Sparse DAG cannot model this ambiguity, as it does not allow multiple expanded states being chosen for a single [MASK] token. This limitation significantly restricts the model’s ability to capture rich semantic variations, leading to a performance drop.
@@
ExLM w/ Sparse DAG 84.4 91.2 92.0 56.9 81.1
Vanila MLM 83.6 90.0 90.3 54.7 79.6
ExLM 85.1 91.8 92.2 57.5 81.6

### v3

@@ I.4 Space Complexity Analysis
However, as shown in Table 13, ExLM with the Sparse DAG experiences some performance degradation. While it still outperforms the version of ExLM without a transition matrix (ExLM w/o Transitions), it falls behind the original ExLM. The reason for this decline is that, in the Sparse DAG, only one expanded state can be selected for each [MASK] token, which prevents the model from capturing longer but possible outputs. For example, in the sentence “This is [MASK], and I’m very [MASK] to see this,” both “so good” and “happy” are reasonable interpretations for the first and second [MASK] tokens, respectively. However, the Sparse DAG cannot model this ambiguity, as it does not allow multiple expanded states being chosen for a single [MASK] token. This limitation significantly restricts the model’s ability to capture rich semantic variations, leading to a performance drop.
@@
ExLM w/ Sparse DAG 84.4 91.2 91.3 56.9 81.0
Vanilla MLM 83.6 90.0 90.3 54.7 79.6
ExLM 85.1 91.4 91.3 57.6 81.4